(6 pts) If the maintenance margin is $2,700, will a margin call result? If so, on...

  1. (6 pts) If the maintenance margin is $2,700, will a margin call result? If so, on which day? How much will Mike be required to deposit into his account?
  1. (8 pts) Suppose a trader sells 5  S&P 500 e – mini contracts for one month out at a value of 2871.68. At the settlement date, the S&P 500 index is at 2999. What is the traders total profit or loss?
  1. (8 pts) Suppose a trader sells 5  S&P 500 e – mini contracts for one month out at a value of 2871.68. At the settlement date, the S&P 500 index is at 2825.3 What is the traders total profit or loss?
Answer #1

For E-Mini Contract Unit =50 (For Margin Call Calculation Initial Margin and price of S&P index each day of holding is required)

Traders Profit/Loss at expiration price of 2999= (2871.68-2999)*50= - $6366

Traders P/L at expiration price of 2825.3= (2871.68-2825.3)*50=+$2319